x/Isl.exe/1o u-lgNslkr7 8P3)H- Solving a word problem using a system of linear equations of the.. Charlie the trainer has two solo workout plans that he offers his dients: Plan A and Plan B. Each client does either one or the other (not both). On Monday there were 3 dlients who did Plan A and 2 who did Plan B. On Tuesday there were 5 dients who did Plan A and 6 who did Plan B. Charlie trained his Monday dients for a total of 3 hours and his Tuesday dients for a total of 7 hours. How long does each of the workout plans last? Length of each Plan A workout: hour(s) 2 DOLL
Answer #1

Solution :

Length of each plan A workout be ti hours

\mathrm{and\:length\:of\:each\:plan\:B\:workout\:be\:}t_{2}\:\mathrm{hours}.

\mathrm{Then,\:on\:Monday,\:\:3\:clients\:who\:did\:plan\:A\:\:and\:\:2\:clients\:who\:did\:\:plan\:B\:for\:a\:total\:of\:3\:hours.}

3t_1+2t_2=3

\mathrm{Then,\:on\:Tuesday,\:\:5\:clients\:who\:did\:plan\:A\:\:and\:\:6\:clients\:who\:did\:\:plan\:B\:for\:a\:total\:of\:7\:hours.}

5t_1+6t_2=7

\mathrm{Following\:are\:the\:two\:systems\:of\:equations\::}

\begin{bmatrix}3t_1+2t_2=3\\ 5t_1+6t_2=7\end{bmatrix}

\mathrm{Using\:the\:substitution\:method\:for\:systems\:of\:equations,}

\mathrm{Isolate}\:t_1\:\mathrm{for}\:3t_1+2t_2=3:

3t_1+2t_2=3

\mathrm{Subtract\:}2t_2\mathrm{\:from\:both\:sides}

3t_1+2t_2-2t_2=3-2t_2

\mathrm{Simplify}

3t_1=3-2t_2

\mathrm{Divide\:both\:sides\:by\:}3

\frac{3t_1}{3}=\frac{3}{3}-\frac{2t_2}{3}

\mathrm{Simplify}

t_1=\frac{3-2t_2}{3}

\mathrm{Substitute\:}t_1=\frac{3-2t_2}{3}

\begin{bmatrix}5* \frac{3-2t_2}{3}+6t_2=7\end{bmatrix}

\mathrm{Isolate}\:t_2\:\mathrm{for}\:5*\frac{3-2t_2}{3}+6t_2=7:

5* \frac{3-2t_2}{3}+6t_2=7

\frac{5\left(3-2t_2\right)}{3}+6t_2=7

\mathrm{Multiply\:both\:sides\:by\:}3

\frac{5\left(3-2t_2\right)}{3}*3+6t_2*3=7*3

\mathrm{Refine}

5\left(3-2t_2\right)+18t_2=21

\mathrm{Add\:similar\:elements:}\:-10t_2+18t_2=8t_2

15+8t_2=21

\mathrm{Subtract\:}15\mathrm{\:from\:both\:sides}

15+8t_2-15=21-15

\mathrm{Simplify}

8t_2=6

\mathrm{Divide\:both\:sides\:by\:}8

\frac{8t_2}{8}=\frac{6}{8}

\mathrm{Simplify}

t_2=\frac{3}{4}

\mathrm{For\:}t_1=\frac{3-2t_2}{3}

\mathrm{Substitute\:}t_2=\frac{3}{4}

t_1=\frac{3-2* \frac{3}{4}}{3}=\frac{3-\frac{3}{2}}{3}=\frac{\frac{3*2-3}{2}}{3}=\frac{\frac{6-3}{2}}{3}=\frac{\frac{3}{2}}{3}=\frac{3}{6}=\frac{1}{2}

\mathrm{Substitute\:}t_{1}=\frac{1}{2}\:\mathrm{\:in\:the\:equation\:\:}3t_1+2t_2=3\::

3*\left ( \frac{1}{2} \right )+2t_2=3

\frac{3}{2}+2t_2=3

\mathrm{Subtract\:}\frac{3}{2}\mathrm{\:from\:both\:sides}

\frac{3}{2}+2t_2-\frac{3}{2}=3-\frac{3}{2}

\mathrm{Simplify}

2t_2=\frac{3}{2}

\mathrm{Divide\:both\:sides\:by\:}2

\frac{2t_2}{2}=\frac{\frac{3}{2}}{2}

\mathrm{Simplify}

t_2=\frac{3}{4}

Hence

\mathrm{Length\:of\:each\:plan\:A\:workout\:=\:\frac{1}{2}\:hour(s).}

\mathrm{and\:length\:of\:each\:plan\:B\:workout\:=\:\frac{3}{4}\:hour(s).}